OpenStudy (anonymous):

Find an equation for the tangent line to the given curve at the point where x=x[0]. Give your answer in Y=mx+b form, where m and b are simplified fractions. y=(x)/(7x+9) ; x[0] =-1 Please help is calculus!!

OpenStudy (anonymous):

hi did you get the derivative of the equation? y=(x)/(7x+9)

OpenStudy (anonymous):

vdu/dx-udv/dx use y' = --------------- v^2

OpenStudy (anonymous):

How do i use it. I am stuck right at this place. (7)/(7x+9)^(2)

OpenStudy (anonymous):

ok that was y'=m=(9)/(7x+9)^(2)]x=-1 solve this one with x=-1 and that will be your m=slope

OpenStudy (anonymous):

Thx, I can get thru that part. But they want the tangent line. In fractions thats why I am confused.

OpenStudy (anonymous):

your m=slope=9/4 then use this m on y-y1=m(x-x1)

OpenStudy (anonymous):

Is x[0]=-1

OpenStudy (anonymous):

ok at x=x(0)=-1 your y(-1)=(-1)/(7(-1)+9) =-1/2.. now youre at position (x1,y1)=(-1,-1/2) use this also including you value of m=9/4, at the eq of line y-y1=m(x-x1)

OpenStudy (anonymous):

now your eq of the line tangent will be y-y1=m(x-x1) y-(-1/2)=(9/4)(x-(-1)) y+1=(9/4)(x+1) eq of the tangent line

OpenStudy (anonymous):

Thx, Not sure why they got 2 1/4x + 1 3/4 as an answer. I am lost.

OpenStudy (anonymous):

or you can put it in the form of Y=mx+b form,

OpenStudy (anonymous):

Y=mx+b y+1=(9/4)(x+1) y=(9/4)(x+1) +1 now solve for y y=9x/4 +9/4 +1 y=9x/4 +13/4 ans

OpenStudy (anonymous):

Thanks so much, i will copy this steps, but the answer was different.

OpenStudy (anonymous):

ok yw,, gud luck now

OpenStudy (anonymous):

I ll keep trying! thx again..